I need to graph a pencil of lines described by y = mx + 3m - 2. Can you help me?

Answer by Nate(3500)

I don't know what you are asking when you say a pencil of lines, but here is what I know:
y = mx + 3m - 2
y = m(x + 3) - 2
y + 2 = m(x + 3)
slope = m
(-3,-2) is a point